# Refuting Perfect Matchings in Spectral Expanders is Hard

### v1

@@ 2.1.1 Affine restriction
An important lemma we will need is that given a set of axioms $\mathcal{Q}$ over the ring $\mathbb{F}[x_{1},\dots,x_{n}]$, a partial assignment of variables can only make refuting $\mathcal{Q}$ easier. Given a set of $m$ polynomial equality constraints $\mathcal{Q}$ over boolean variables $\{x_{1},\dots,x_{n}\}$, let the family of functions $\{f_{i}:\{0,1\}^{n}\rightarrow\{\texttt{True},\texttt{False}\}\}_{i\in[m]}$, denote predicates for satisfiability for each constraint. For example, given $\alpha\in\{0,1\}^{n}$, $f_{i}(\alpha)=\texttt{True}$ if the $i$’th polynomial constraint $q_{i}\in\mathcal{Q}$ is satisfied i.e $p_{i}(\alpha)=0$. We say $\mathcal{Q}$ is satisfied if $\exists\alpha\in\{0,1\}^{n}$ such that $f_{i}(\alpha)=\texttt{True}\iff q_{i}(\alpha)=0$ for all $i\in[m]$. Given a map $\rho:\{x_{1},\dots,x_{n}\}\rightarrow\{x_{1},\dots,x_{n},\bar{x}_{1},\dots,\bar{x}_{n},1,0\}$, the restriction of a function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, denoted by $f|_{\rho}$, is defined as $f|_{\rho}(x_{1},\dots,x_{n})=f(\rho(x_{1}),\dots,\rho(x_{n}))$. Similarly, the restriction of formula $\mathcal{Q}$ is defined as $\mathcal{Q}|_{\rho}=\{f_{1}|_{\rho},\dots,f_{m}|_{\rho}\}$. Two formula $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are equivalent if they are element-wise equal, ignoring any functions that are constantly True. For example, $\mathcal{Q}=\{f_{a},f_{b},\texttt{True}\}$ and $\mathcal{Q}^{\prime}=\{f_{a},f_{b}\}$ are equivalent, denoted as $\mathcal{Q}\equiv\mathcal{Q}^{\prime}$.
@@
The proof for the above lemma can be found in [Austrin_2022, Lemma 2.2]. What the above lemma says is that if we have a graph $G$ with odd vertices woth constant degree, that has a perfect matching on a subset of even vertices on the graph, then the size of the proof to refute $\text{PM}(G)$ is at least as large as refuting a perfect matching in $G$ with the even vertices removed.
@@ 2.2 Graph Theory Preliminaries
If a graph $G$ has even number of vertices and for every subset $S\subseteq G$ we have $q(G\setminus S)\leq|S|$, then $G$ contains a perfect matching.
@@
$q(G\setminus(S\cup T))\leq|S|f-\sum_{w\in T}(f-|\Gamma_{G}\left(w\right)\setminus S|).$
@@ 2.3 Probabilistic Tools
Suppose $X_{1},...,X_{n}$ are identical independent random variables taking values in $\{0,1\}$. Let $X$ denote their sum and let $\mu=n\mathbb{E}\left[X_{1}\right]$ denote the sum’s expected value. Then, for any $10<\delta<$ 1, we have
@@
    For all $i\in[n]$, $\Pr_{\mathcal{D}}\left[E_{i}\right]\leq\beta$
@@
[…]
The proof of (Multiplicative Chernoff bound) and (Lovàsz Local Lemma) can be found in any textbook on randomised algorithms (for example, see [mitzenmacher2017probability, Chapter 1, Chapter 7]). The following is the main result of this section.
@@
For every $10<c<$ 1 and $\gamma>0$, there exists $d_{0}$ such that the following holds. If $G$ is a $d$-regular graph, for some $d\geq d_{0}$, then there exists a subset $A\subseteq V(G)$ such that
@@ 3 Topological embedding
$\left|2m-\frac{d}{n}b^{2}\right|\leq\lambda b.$
Dividing both sides with $b(b-1)$, and observing that $\frac{db^{2}}{n(b-1)b}=\frac{d}{n}\frac{b}{b-1}=\frac{d}{n}(1+\frac{1}{b-1})$, we further get
@@
$\leq\frac{2\lambda}{b}.$ (6)
Let us briefly justify each step: (4) comes from the triangle inquality and $d/n\leq 1$; Equation (5) comes from Equation (2); the last inequality comes from the fact that $b\in\Omega\left(n\right)$, so the inequality holds for $n$ large enough.
Let $\xi$ be as given by (Embedding Theorem). Using the bound on the difference between $p$ and $d/n$, for every disjoint subsets $S,T\subseteq B$ of size $|S|,|T|\geq\xi n$, for $\lambda<\varepsilon d$ where $\varepsilon$ is sufficiently small, we have
$\Bigl{|}e_{G}(S,T)-p|S||T|\Bigr{|}$ $\leq\left|e_{G}(S,T)-\frac{d}{n}|S||T|\right|+\left|\frac{d}{n}|S||T|-p|S||T|\right|$
$\leq\lambda\sqrt{|S||T|}+\frac{2\lambda}{b}|S||T|\leq\xi|S||T|p.$
With the lower bounds on $S,T$ and $B$, we can make $\varepsilon$ sufficiently small with respect to $\xi$ to get the upper bound in the last step. Let $\sigma:E(H)\to\mathbb{N}$ be the constant function where $\sigma(e)$ is the smallest odd integer larger than $\log n$. As $G[B]$ has at least $Cn$ edges (by (Expander Mixing Lemma)), $\sigma(e)\leq 2+\log n$, and $H$ has at most $\alpha\frac{n}{\log n}$ vertices, we can invoke (Embedding Theorem) to conclude that $G[B]$ contains $H^{\sigma}$. ∎
@@ 4 Perfect matching and regular subgraphs
As described earlier, the second ingredient in our hardness proof is showing that a certain residual graph contains a perfect matching or a spanning $(t-1)$-regular subgraph. In this section we state and prove these ingredients.
@@
Suppose $G$ is an $(n,d,\lambda)$-graph, where $\lambda<d/50$. Suppose $U\subseteq V(G)$ is such that $\delta(G[U])\geq\frac{9d}{10}$. Then for all $S\subseteq U$, $G[U\setminus S]$ has at most $|S|$ connected components, that is, $q(G[U\setminus S])\leq|S|$.
@@
Consider some $U\subseteq V(G)$ such that $\delta(G[U])\geq\frac{9d}{10}$. Consider some $S\subseteq U$. We aim to show that $G[U\setminus S]$ has at most $|S|$ connected components. If $|S|\geq|U|/2$ then as $G[U\setminus S]$ has fewer than $|S|$ vertices, so the the upper bound on connected components trivially holds. Thus, we can assume $|S|<|U|/2$. We claim the following:
@@
If $\text{If }e_{G}(X,Y)\geq 1\text{ for every partition $X\cup Y=U\setminus S$, $|X|,|Y|\geq|S|/3$, then }q(G[U\setminus S])\leq|S|.$ (7)
@@
To see why, assume towards a contradiction that there exists an edge between every partition $X\cup Y=U\setminus S$, where $|X|,|Y|\geq|S|/3$, and $G[U\setminus S]$ has more than $|S|$ connected components. Denote the vertex sets of these components by $C_{1},\ldots,C_{k}$, for some $k>|S|$. Let $X^{\star}:=C_{1}\cup\ldots\cup C_{s}$ and $Y^{\star}:=C_{s+1}\cup\ldots\cup C_{k}$, where $s=\lfloor|S|/2\rfloor\geq|S|/3$. By construction, even if each component $C_{i}$ is a singleton set, we get that $|X^{\star}|,|Y^{\star}|\geq|S|/3$. Now as all $C_{i}$’s are disjoint connected components, there can be no edge between $X^{\star}$ and $Y^{\star}$. Therefore, we have found a correct sized partition of $U\setminus S$ without an edge between them, which contradicts our assumption.

### v2

@@ 2.1.1 Affine restriction
An important lemma we will need is that given a set of axioms $\mathcal{Q}$ over the ring $\mathbb{F}[x_{1},\dots,x_{n}]$, a partial assignment of variables can only make refuting $\mathcal{Q}$ easier. Given a set of $m$ polynomial equality constraints $\mathcal{Q}$ over boolean variables $\{x_{1},\dots,x_{n}\}$, let the family of functions $\{f_{i}:\{0,1\}^{n}\rightarrow\{\texttt{True},\texttt{False}\}\}_{i\in[m]}$, denote predicates for satisfiability for each constraint. For example, given $\alpha\in\{0,1\}^{n}$, $f_{i}(\alpha)=\texttt{True}$ if the $i$’th polynomial constraint $q_{i}\in\mathcal{Q}$ is satisfied i.e $q_{i}(\alpha)=0$. We say $\mathcal{Q}$ is satisfied if $\exists\alpha\in\{0,1\}^{n}$ such that $f_{i}(\alpha)=\texttt{True}\iff q_{i}(\alpha)=0$ for all $i\in[m]$. Given a map $\rho:\{x_{1},\dots,x_{n}\}\rightarrow\{x_{1},\dots,x_{n},\bar{x}_{1},\dots,\bar{x}_{n},1,0\}$, the restriction of a function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, denoted by $f|_{\rho}$, is defined as $f|_{\rho}(x_{1},\dots,x_{n})=f(\rho(x_{1}),\dots,\rho(x_{n}))$. Similarly, the restriction of formula $\mathcal{Q}$ is defined as $\mathcal{Q}|_{\rho}=\{f_{1}|_{\rho},\dots,f_{m}|_{\rho}\}$. Two formula $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are equivalent if they are element-wise equal, ignoring any functions that are constantly True. For example, $\mathcal{Q}=\{f_{a},f_{b},\texttt{True}\}$ and $\mathcal{Q}^{\prime}=\{f_{a},f_{b}\}$ are equivalent, denoted as $\mathcal{Q}\equiv\mathcal{Q}^{\prime}$.
@@
The proof for the above lemma can be found in [Austrin_2022, Lemma 2.2]. What the above lemma says is that if we have a graph $G$ with odd vertices with constant degree, that has a perfect matching on a subset of even vertices on the graph, then the size of the proof to refute $\text{PM}(G)$ is at least as large as refuting a perfect matching in $G$ with the even vertices removed.
@@ 2.2 Graph Theory Preliminaries
If a graph $G$ has even number of vertices and for every subset $S\subseteq V(G)$ we have $q(G\setminus S)\leq|S|$, then $G$ contains a perfect matching.
@@
$q(G\setminus(S\cup T))\leq|S|f-\sum_{w\in T}(f-|\Gamma_{G}\left(w\right)\setminus S|)$
@@ 2.3 Probabilistic Tools
Suppose $X_{1},...,X_{n}$ are identical independent random variables taking values in $\{0,1\}$. Let $X$ denote their sum and let $\mu=n\mathbb{E}\left[X_{1}\right]$ denote the sum’s expected value. Then, for any $0<\delta<1$, we have
@@
    For all $i\in[n]$, $\Pr_{\mathcal{D}}\left[E_{i}\right]\leq\beta.$
@@
$\Pr{\overset{n}{\underset{i=1}{\bigcap}}\hskip 2.84544pt\overline{E_{i}}}{\mathcal{D}}>0$
The proof of (Multiplicative Chernoff bound) and (Lovàsz Local Lemma) can be found in any textbook on randomised algorithms (for example, see [mitzenmacher2017probability, Chapter 1, Chapter 7]). The following lemma is by [Austrin_2022, Lemma 4.3], re-derived here for completeness.
@@
For every $0<c<1$ and $\gamma>0$, there exists $d_{0}$ such that the following holds. If $G$ is a $d$-regular graph, for some $d\geq d_{0}$, then there exists a subset $A\subseteq V(G)$ such that
@@ 3 Topological embedding
$\left|2m-\frac{d}{n}b^{2}\right|\leq\lambda b$
Dividing both sides with b(b1)b(b-1), and observing that $\frac{db^{2}}{n(b-1)b}=\frac{d}{n}\frac{b}{b-1}=\frac{d}{n}(1+\frac{1}{b-1})$, we further get
@@
$\leq\frac{2\lambda}{b}$ (5)
Let us briefly justify each step: (3) comes from the triangle inquality and $d/n\leq 1$; Equation (4) comes from Equation (1); the last inequality comes from the assumption that $b\in\Omega\left(n\right)$, so the inequality holds for $n$ large enough.
Let $\xi$ be as given by the (Embedding Theorem). Using the bound on the difference between $p$ and $d/n$, for every disjoint subsets $S,T\subseteq B$ of size $|S|,|T|\geq\xi n$, for $\lambda<\varepsilon d$ where $\varepsilon$ is sufficiently small, we have
[…]
$\leq\lambda\sqrt{|S||T|}+\frac{2\lambda}{b}|S||T|\leq\xi|S||T|p$
With the lower bounds on $S,T$ and $B$, we can make $\varepsilon$ sufficiently small with respect to $\xi$ to get the upper bound in the last step. Let $\sigma:E(H)\to\mathbb{N}$ be the constant function where $\sigma(e)$ is the smallest odd integer larger than $\log n$. As $G[B]$ has at least $Cn$ edges (by the (Expander Mixing Lemma)), $\sigma(e)\leq 2+\log n$, and $H$ has at most $\alpha\frac{n}{\log n}$ vertices, we can invoke the (Embedding Theorem) to conclude that $G[B]$ contains $H^{\sigma}$. ∎
@@ 4 Perfect matching and regular subgraphs
As described earlier, the second ingredient in our hardness proof is showing that a certain residual graph contains a perfect matching or a spanning (t1)(t-1)-regular subgraph. In this section we state and prove these ingredients.
@@
Let $G$ be an $(n,d,\lambda)$-graph with $\lambda<d/50$, and suppose $G^{\prime}\subseteq G$ satisfies $\delta(G^{\prime})\geq 0.9d$. Then for all $S\subseteq V(G^{\prime})$, $G^{\prime}\setminus S$ has at most $|S|$ connected components, that is, $q(G^{\prime}\setminus S)\leq|S|$. Therefore, if $G^{\prime}$ has an even number of vertices then it contains a perfect matching.
@@
Let $U=V(G^{\prime})$. We aim to show that the graph $G^{\prime}\setminus S$ has at most $|S|$ connected components. If $|S|\geq|U|/2$ then $G^{\prime}\setminus S$ has at most $|S|$ vertices, so the the upper bound on connected components trivially holds. For the remainder of the proof we can assume $|S|<|U|/2$.
We claim the following:
@@
$\text{For every partition }X\cup Y=U\setminus S,\text{ with }|X|,|Y|\geq|S|/3,\qquad e_{G^{\prime}}(X,Y)\geq 1\implies q(G^{\prime}\setminus S)\leq|S|.$ (6)
@@
To see why, assume towards a contradiction that there exists an edge in $G^{\prime}$ between every partition $X\cup Y=U\setminus S$, where $|X|,|Y|\geq|S|/3$, and $G^{\prime}\setminus S$ has more than $|S|$ connected components. Denote the vertex sets of these components by $C_{1},\ldots,C_{k}$, for some $k>|S|$. Let $X^{\star}:=C_{1}\cup\ldots\cup C_{s}$ and $Y^{\star}:=C_{s+1}\cup\ldots\cup C_{k}$, where $s=\lfloor|S|/2\rfloor\geq|S|/3$. By construction, even if each component $C_{i}$ is a singleton set, we get that $|X^{\star}|,|Y^{\star}|\geq|S|/3$. Now as all $C_{i}$’s are disjoint connected components, there can be no edge between $X^{\star}$ and $Y^{\star}$. Therefore, we have found a partition $X^{\star}\cup Y^{\star}=U\setminus S$ with $|X^{\star}|,|Y^{\star}|\geq|S|/3$ without an edge between them, which contradicts our assumption that all appropriately sized partitions have at least one edge between them.
To complete our proof, it suffices to show $e_{G^{\prime}}(X,Y)\geq 1$ for every partition $X\cup Y=U\setminus S$ with $|X|,|Y|\geq|S|/3$.
Consider some arbitrary partition $X\cup Y$ of $U\setminus S$, with $|X|,|Y|\geq|S|/3$, and without loss of generality assume $|X|\leq|Y|$. Then by a simple counting argument we get
$|X|\leq(|U|-|S|)/2\leq(n-|S|)/2$ (7)
We have:
$e_{G^{\prime}}(X,X)+e_{G^{\prime}}(X,S)$ $\leq e_{G}(X,X)+e_{G}(X,S)$
$\leq\frac{d}{n}|X|^{2}+\lambda|X|+\frac{d}{n}|X||S|+\lambda\sqrt{|X||S|}$ (8)
$\leq\frac{d}{n}|X|\frac{(n-|S|)}{2}+\lambda|X|+\frac{d}{n}|X||S|+\lambda\sqrt{3}|X|$ (9)
$<\frac{d|X|}{2}+\frac{d|X||S|}{2n}+3\lambda|X|$ (10)
$<\frac{d}{2}|X|+\frac{d}{4}|X|+3\lambda|X|$ (11)
$<\frac{9d}{10}|X|$ (12)
These steps are justified as follows: The first equation follows from the (Expander Mixing Lemma); Equation (9) comes from Equation (7) and $|X|\geq|S|/3$; Equation (11) comes from $|S|<\frac{n}{2}$; and Equation (12) comes the assumption $\lambda<d/50$. By the assumption $\delta(G^{\prime})>0.9d$ we conclude that there is an edge in $G^{\prime}$ with one vertex in $X$ and the other in $V(G^{\prime})\setminus(X\cup S)=Y$. ∎
The next lemma shows that subgraphs of $(n,d,\lambda)$-graphs with large minimum degree contain regular spanning subgraphs.
Lemma 4.2 (Regular Subgraph Lemma)
For every $C>1$ there exists $d_{0}=d_{0}(C)$ such that the following holds. Suppose $G$ is an $(n,d,\lambda)$ graph with $\lambda<\varepsilon d$ and $d\geq d_{0}$, where $\varepsilon<1/100C^{3/2}$. If $G^{\prime}\subseteq G$ has minimum degree $\delta({G^{\prime}})\geq d-C$, then $G^{\prime}$ contains a spanning $f$-regular subgraph for any even $2\leq f\leq d/2$.
Proof.
We prove this lemma using the (Tutte’s Generalised Criterion). We need to show that for any pair of disjoint sets $S,T\subseteq V(G^{\prime})$, we have
$q(G^{\prime}\setminus(S\cup T))$ $\leq|S|f-\sum_{w\in T}(f-|\Gamma_{G^{\prime}}\left(w\right)\setminus S|)$ (13)
As $\varepsilon<\frac{1}{100C^{3/2}}$ and $C>1$, we have that $\varepsilon<1/100$. This implies that $G$ is an $(n,d,\lambda)$ graph with $\lambda<d/100$. We set $d_{0}:=d_{0}(C)$ large enough such that for all $d\geq d_{0}$, even after deleting at most $C$ edges incident on each vertex of the $d$-regular graph $G$ to get $G^{\prime}$, we have the minimum degree of $G^{\prime}$ to be $\delta({G^{\prime}})\geq d-C>9d/10$. Therefore the conditions of Lemma 4.1 are satisfied, thus
$q(G^{\prime}\setminus(S\cup T))\leq|S\cup T|=|S|+|T|$
To prove (13), it suffices to show
$|S|+|T|$ $\leq|S|f-|T|f+|T|(d-C)-e_{G^{\prime}}(S,T)$ (14)
$\leq|S|f-\sum_{w\in T}(f-|\Gamma_{G^{\prime}}\left(w\right)\setminus S|)$
We distinguish a few cases.
Case 1:
Suppose $|S|\leq|T|$. As $f\leq d/2$, we have
$|S|f+|T|(d-C-f)$ $\geq(|S|+|T|)(d/2-C)$ (15)
The condition described by the inequality (14) is satisfied via the following analysis.
$|S|+|T|+e_{G^{\prime}}(S,T)$ $\leq|S|+|T|+|S||T|\frac{d}{n}+\varepsilon d\sqrt{|S||T|}$ (16)
$\leq|S|+|T|+\frac{d}{4}(|S|+|T|)+\varepsilon d\frac{1}{2}(|S|+|T|)$ (17)
$\leq(|S|+|T|)\left(\frac{d}{4}+1+\frac{\varepsilon d}{2}\right)$ (18)
$<(|S|+|T|)\left(d/2-C\right)$ (19)
$\leq|S|f-|T|(d-C-f)$ (20)
Equation (16) comes from the (Expander Mixing Lemma) and $\lambda<\varepsilon d$, together with an obvious upper bound $e_{G^{\prime}}(S,T)\leq e_{G}(S,T)$. Equation (17) comes from the fact that $|S||T|\leq\left(\frac{|S|+|T|}{2}\right)^{2}\leq n\frac{(|S|+|T|)}{4}$. Equation (19) comes from $\varepsilon<\nicefrac{{1}}{{100C^{3/2}}}$, $C>1$ and $d_{0}$ being sufficiently large. Equation (20) follows from Equation (15) which gives us what we want.
Case 2:
Suppose $|S|>|T|$. As $f\geq 2$, we have
$|S|f+|T|(d-C-f)$ $\geq 2|S|+|T|(d-C-2)$ (21)
To show Equation (14), it suffices to show that
$e_{G^{\prime}}(S,T)\leq|S|+|T|(d-C-3)$ (22)
Now we distinguish between two subcases.
  1. 1.
    If $|T|\leq\frac{|S|}{C+3}$, then Equation (2) follows from a trivial bound $e_{G^{\prime}}(S,T)\leq|T|d$.
  2. 2.
    $\frac{|S|}{C+3}<|T|<|S|$. As $|S|+|T|<n$ we have $|S|<n-\frac{|S|}{C+3}$, thus $|S|<n\left(\frac{C+3}{C+4}\right)$. Using the (Expander Mixing Lemma), we have
    $e_{G}(S,T)\leq e_{G^{\prime}}(S,T)$ $\leq\frac{d}{n}|S||T|+\varepsilon d\sqrt{|S||T|}$
    $<d\frac{(C+3)}{(C+4)}|T|+\varepsilon d|T|\sqrt{(C+3)}$
    $=|T|d\left(\frac{C+3}{C+4}+\varepsilon\sqrt{C+3}\right)$
    $<|T|d\left(\frac{C+3}{C+4}+\frac{1}{2(C+4)}\right)=|T|d\left(1-\frac{1}{2(C+4)}\right)$
    where the penultimate inequality follows from the upper bound on $\varepsilon$. For $d$ sufficiently large in terms of $C$ we obviously have
    $|T|d\left(1-\frac{1}{2(C+4)}\right)<|T|d-|T|(C+3)<|S|+|T|(d-C-3)$
    hence is satisfied.
5 Proof of Theorem 1.1
In this section we prove Theorem 1.1. As $\text{Card}(G,\mathbf{t})\equiv\text{Card}(G,\mathbf{d-t})$, without loss of generality we only prove the theorem for $t\leq d/2$.
Proof of Theorem 1.1.
Let $G=(V,E)$ be an $(n,d,\lambda)$-graph on an odd number of vertices with $\lambda<\varepsilon d$, where $\varepsilon<\nicefrac{{1}}{{100C^{3/2}}}$ and $C=6$. For sufficiently small constant $\alpha\in(0,1)$, let $H$ denote the graph on $h=\alpha\frac{n}{\log n}$ vertices as given by Lemma 2.5 (to show lower bounds for PC, we use $H$ from Lemma 2.4). Recall that any SoS proof which refutes $\text{PM}(H)$ has degree $\Omega\left(h\right)$. We now make use of $H$ to show the hardness of refuting $\text{Card}(G,\mathbf{t})$. The idea is to find a restriction $\rho$ such that $\text{Card}(G,\mathbf{t})|_{\rho}\equiv\text{PM}(H)$. We achieve this through the following steps.
  1. 1.
    Invoke Lemma 2.14 with parameters $c=0.925$ and $\gamma=0.025$ to get subsets $A\subseteq V(G)$ and $B=V(G)\setminus A$, such that for every $u\in V(G)$ we have
    $0.9d$ $\leq|\Gamma_{G}\left(u\right)\cap A|\leq 0.95d$ (23)
    $0.05d$ $\leq|\Gamma_{G}\left(u\right)\cap B|\leq 0.1d$ (24)
  2. 2.
    From equations (23) and (24), $|B|>n/20$, with room to spare and $|E(G[B])|\geq nd/20$. By Corollary 3.3, $G[B]$ contains $H^{\sigma}$ such that each $\sigma(e)$ is odd. Let us denote a subgraph of $G[B]$ corresponding to $H^{\sigma}$ by $G_{\Psi}$. We can describe $G_{\Psi}$ as a function $\Psi:V(H)\to B$ together with a collection of pairwise vertex-disjoint (other than at the endpoints) paths $\mathsf{P}_{\Psi\left(u\right)\rightsquigarrow\Psi\left(v\right)}$ in $G[B]$, for $(u,v)\in E(H)$. Observe that it is at least as hard to refute [^5] Note that this by itself does not guarantee it is hard to refute $\text{PM}(G)$. We need item (3) and this to show hardness of refuting $\text{PM}(G)$. $\text{PM}(G_{\Psi})$ as it is to refute $\text{PM}(H)$. To see why, let $y_{1},\dots,y_{E(H)}$ denote the variables for the $\text{PM}(H)$ formulae for each edge of $H$. We use as shorthand $\mathcal{Y}=(y_{e})_{e\in E(H)}$ and $\overline{\mathcal{Y}}=(\overline{y}_{e})_{e\in E(H)}$. Define a mapping $\rho^{\prime}:E(G_{\Psi})\to\{0,1,\mathcal{Y},\overline{\mathcal{Y}}\}$ as follows. For each $(u,v)\in E(H)$, let $\rho^{\prime}(x_{e})=y_{u,v}$ where $e$ is the first edge on the path $\mathsf{P}_{\Psi\left(u\right)\rightsquigarrow\Psi\left(v\right)}$. Map each variable $x_{e}$ for $e\in\mathsf{P}_{\Psi\left(u\right)\rightsquigarrow\Psi\left(v\right)}$ alternately to $y_{u,v}$ or $\bar{y}_{u,v}$, such that the edges of the path adjacent to $\Psi\left(u\right)$ and $\Psi\left(v\right)$ are set to $y_{u,v}$. This is always possible as $\mathsf{P}_{u\rightsquigarrow v}$ has odd length. Observe that $\text{PM}(G_{\Psi})|_{\rho^{\prime}}\equiv\text{PM}(H)$.
  3. 3.
    As $n$ is odd and $|V(G_{\Psi})|$ is odd, we have that $U=V(G)\setminus V(G_{\Psi})$ has even size. From (23) we have that $G[U]$ has minimum degree at least $\frac{9d}{10}$. As $\lambda<d/50$ (with room to spare), we can invoke Lemma 4.1 to conclude $G[U]$ has a perfect matching $M$.
  4. 4.
    Consider the subgraph $G^{\prime}\subseteq G$ obtained by deleting all edges $e\in E(G_{\Psi})\cup M$, where $M$ is the perfect matching from the step above. As $\Delta_{H}\leq 5$, every vertex $u\in G$ loses at most $5$ edges in this process. Thus, we have $\delta({G^{\prime}})\geq d-5$. As $\lambda<\varepsilon d$, by the (Regular Subgraph Lemma) we have that $G^{\prime}$ contains a (t1)(t-1)-regular spanning subgraph $G^{\prime\prime}$.
We finally define $\rho$ as follows:
$\rho(x_{e})=\left\{\begin{array}[]{@{}cl}\rho^{\prime}(e)&\text{if }e\in E(G_{\Psi})\\ 1&\text{if }e\in M\cup E(G^{\prime\prime})\\ 0&\text{otherwise.}\end{array}\right$.
Then $\text{Card}(G,\mathbf{t})|_{\rho}\equiv\text{PM}(H)$, thus our theorem follows from Lemma 2.7. ∎
6 Related Work And Concluding Remarks
In propositional [^6] As opposed to algebraic proof complexity proof complexity, there a few prior examples of the strategy of embedding a worst case instance into a host graph to show lower bounds for a larger class of objects [itsykson2021Near, pitassi2016PolyLogFrege]. Pitassi, Rossman, Servedio and Tan show Tseitin lower bounds for Frege proof systems [pitassi2016PolyLogFrege] by relying on the embedding result by Rubinfeld and Kleinberg [kleinberg1996short], which allows one to embed any bounded degree graph $H$ of size $O\left(n/\mathsf{poly}(\log n)\right)$ into an expander graph on $n$ vertices as a minor (not necessarily a topological one). Krivelevich and Nenadov simplify and improve the above embedding theorem to allow for embedding any graph $H$ with size $O\left(n/\log n\right)$ as an ordinary minor [krivelevich2021completeMinors]. However, embedding a hard instance $H$ into $G$ as an ordinary minor does not guarantee that the hardness of $H$ is preserved in the setting considered in this paper. In particular, it is entirely possible that one of the edge contractions to obtain the minor results in $H$ now being easy to refute. Thus, these embedding theorems cannot be directly applied. Instead, as described in section 5, one way to preserve hardness is to use embedding theorems that allow for topological embeddings that allow for edge sub-divisions of odd size [draganic22rolling, nenadov2023routing]. In order to get a topological embedding, Austrin and Risse modify the ordinary embedding theorem in [krivelevich2021completeMinors] but critically rely on the host graph being random. In this work, we use the embedding theorem by Draganić, Krivelevich, and Nenadov [draganic22rolling], which greatly simplifies the argument. Moreover, we avoid the use of the contiguity argument present in [Austrin_2022] by directly utilising Tutte’s criterion and the Expander Mixing Lemma.
In summary, we show degree lower bounds for refuting $\text{Card}(G,\mathbf{t})$ for odd $t$ in $(n,d,\lambda)$ graphs in the SoS and PC proof systems. There is still a $\log n$ gap between the largest possible proof in such systems, and our lower bounds (similar to Austrin and Risse). It is not inherently clear that such a gap should exist. The gap is an artefact of $d$ being constant, which makes the graphs sparse i.e we need $\Theta(\log n)$ edges to form a path between any two nodes. This implies, that $\Omega\left(n/\log n\right)$ is the largest hard instance we can topologically embed in any graph. Thus, if the worst case lower for refuting perfect matchings was indeed $\Omega(n)$, we would need a more direct proof of the statement without using a smaller hard instance. We leave the issue of resolving the tightness of our lower bound as an open problem for future work.